OpenStudy (anonymous):

The sum of two consecutive integers is 27

OpenStudy (anonymous):

13 14

jimthompson5910 (jim_thompson5910):

let x be the first integer so that makes x+1 be the next consecutive integer

jimthompson5910 (jim_thompson5910):

add them to get x + (x+1) = 2x + 1 then you set that result equal to 27 and solve for x 2x + 1 = 27 2x = 27 - 1 2x = 26 x = 26/2 x = 13 The first (smaller) number is 13 x+1 = 13 +1 x+1 = 14 and the next (larger) number is 14

jimthompson5910 (jim_thompson5910):

Sure enough, 13+14 = 27 and our answer is confirmed